# APOLLO: Automated LLM and Lean Collaboration for Advanced Formal Reasoning

### v3

@@ Abstract
Formal reasoning and automated theorem proving constitute a challenging subfield of machine learning, in which machines are tasked with proving mathematical theorems using formal languages like Lean. A formal verification system can check whether a formal proof is correct or not almost instantaneously, but generating a completely correct formal proof with large language models (LLMs) remains a formidable task. The usual approach in the literature is to prompt the LLM many times (up to several thousands) until one of the generated proofs passes the verification system. In this work, we present APOLLO (Automated PrOof repair via LLM and Lean cOllaboration), a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with an LLM’s reasoning abilities to achieve better proof-generation results at a low sampling budget. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The repaired sub-proofs are recombined and reverified, iterating up to a user-controlled maximum number of attempts. On the miniF2F benchmark, we establish a new state-of-the-art accuracy of 84.9% among sub 8B-parameter models while keeping the sampling budget below one hundred. Moreover, Apollo raises the state-of-the-art accuracy for Goedel-Prover-SFT to 65.6% while cutting sample complexity from 25,600 to a few hundred. General-purpose models (o3-mini, o4-mini) jump from 3–7% to over 40% accuracy. Our results demonstrate that targeted, compiler-guided repair of LLM outputs yields dramatic gains in both efficiency and correctness, suggesting a general paradigm for scalable automated theorem proving. The codebase is available at https://github.com/aziksh-ospanov/APOLLO
@@ 1 Introduction
Formal reasoning has emerged as one of the most challenging fields of AI with recent achievements such as AlphaProof and AlphaGeometry doing well at the International Math Olympiad competing with humans [1, 2, 3]. Formal reasoning relies both on AI models and a formal verification system that can automatically verify whether a mathematical proof is correct or not. In recent years, formal verification systems such as Lean [4] have facilitated a new form of doing mathematical research by enabling mathematicians to interact with the formal verification system and also with each other via the system, enabling larger numbers of mathematicians to collaborate with each other on a single project. As such, these formal verification systems are also called proof assistants as one can use them interactively to write a formal proof and instantaneously observe the current state of the proof and any possible errors or shortcomings in the proof generated by the compiler. Immediate access to the output of Lean compiler can help a mathematician to fix possible errors in the proof. At the same time, when a proof passes the compiler with no error, other collaborators do not need to verify the correctness of the proof. This type of collaboration is transforming the field of mathematics enabling large groups of collaborators to engage in large projects of mathematical research such as the Prime Number Theorem And More project [5]. Moreover, it has given rise to the digitization of mathematics [6].
In the AI front, large language models (LLMs) are improving at mathematical reasoning in natural language, and at the same time, they have shown a remarkable ability to learn the language of those formal verification systems and write mathematical proofs in formal language. This has led to the field of Automated Theorem Proving (ATP) where the standard approach is to prompt the LLM to generate a number of candidate proofs for a given theorem which will then be verified automatically using the formal verification system. Better models, better training sets, and better training methods has led to significant advances in this field [7, 8, 9, 10, 11].
@@
In this work, we address this issue by introducing Apollo, a fully automated system which uses a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with the reasoning abilities of LLMs. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The high-level overview of Apollo appears at the bottom of Figure 1.
@@
  • We establish a new SOTA on miniF2F-test benchmark for medium-sized language models. (with parameter size 7B or less)
@@ 2 Related Works
Formal theorem proving systems. At their core, formal theorem-proving (FTP) systems employ interactive theorem provers (ITPs) to verify mathematical results. In particular, Lean4 [4] is both a functional programming language and an ITP. Each theorem, lemma, conjecture, or definition must be checked by Lean’s trusted kernel, so the compiler’s verdict is binary: either the statement type-checks (True) or it fails (False). This rigorous verification dramatically increases the reliability of formal mathematics; however, it also increases the complexity of proof development: authors must both comprehend the mathematical concepts and precisely encode them in the formal language. The usefulness of Lean is also shown to go beyond theorem proving, as Jiang et al. [12] showed that Lean can be adapted to natural language logical reasoning.
@@
Patching broken programs. Many prior works have explored using feedback to repair broken proofs. In software engineering, “repair” typically refers to program repair, i.e. fixing code that no longer runs [13]. A common source of errors is a version mismatch that introduces bugs and prevents execution. For example, SED [14] uses a neural program-generation pipeline that iteratively repairs initial generation attempts. Other notable approaches train specialized models to fix broken code based on compiler feedback (e.g. BIFI [15], TFix [16]) or leverage unsupervised learning over large bug corpora (e.g. BugLab [17]).
@@
LLM collaboration with external experts. The use of expert feedback, whether from human users or from a proof assistant’s compiler, has proven effective for repairing broken proofs. Ringer [18] showcased automatic proof repair within the Coq proof assistant [19], enabling the correction of proofs in response to changes in underlying definitions. Jiang et al. [20] showed that leveraging automatic solvers with generative models yield better performance on formal math benchmarks. More recently, First et al. [21] showed that incorporating compiler feedback into LLM proof generation significantly improves the model’s ability to correct errors and produce valid proofs. Another line of work [22] explored an idea of mathematician and LLM collaboration, where human experts are aided by LLMs at proof writing stage.
@@
Proof search methods. Numerous recent systems use machine learning to guide the search for formal proofs. One of the methods involves using large language models with structured search strategies, e.g. Best-First Search (BFS) [23, 24, 25] or Monte Carlo Tree Search (MCTS) [26, 27, 2]. While tree-search methods reliably steer a model toward valid proofs, they result in high inference costs and often explore many suboptimal paths before success. Another line of work leverages retrieval based systems to provide context for proof generators with potentially useful lemmas. One such example is ReProver [28] that augments Lean proof generation by retrieving relevant lemmas from a proof corpus and feeding them to an LLM, enabling the model to reuse existing proofs.
@@
Whole proof generation. The use of standalone LLMs for theorem proving has emerged as a major research direction in automated theorem proving. One of the earliest works presented GPT-f [29], a transformer-based model for theorem proving that established the LLMs ability in formal reasoning. As training frameworks and methods advance, the community has produced numerous models that generate complete proofs without external search or tree-search algorithms. Recent work shows that both supervised models [30, 31] and those trained via reinforcement learning [30, 32, 33, 34] achieve competitive performance on formal-math benchmarks. However, whole-proof generation remains vulnerable to hallucinations that can cause compilation failures even for proofs with correct reasoning chains.
@@
Informal Chain-of-Thought in Formal Mathematics. Several recent works demonstrate that interleaving informal “chain-of-thought” (CoT) reasoning with formal proof steps substantially improves whole-proof generation. For example, developers insert natural-language mathematical commentary between tactics, yielding large accuracy gains [31, 30, 35]. Wang, et al. [34] further show that special “thinking” tokens let the model plan its strategy and self-verify intermediate results. The “Draft, Sketch, and Prove” (DSP) [36] and LEGO-Prover [37] pipelines rely on an informal proof sketch before attempting to generate the formal proof. Collectively, these studies indicate that providing informal mathematical context, whether as comments, sketches, or dedicated tokens, guides LLMs toward correct proofs and significantly boosts proof synthesis accuracy.
@@ 3 Our Approach
Our compiler of choice is the Lean REPL [38], a “Read–Eval–Print–Loop” for Lean4. The REPL provides a complete list of compilation errors and warnings, each with a detailed description and the location (for example, the line number) where it occurred. Errors can include syntax mistakes, nonexistent tactics, unused tactics, and more. We aggregate this information to drive our agent’s dynamic code repair and sub-problem extraction. The pseudo-code for our framework can be found in Algorithm 1 in the Appendix.
@@ 3.1 Syntax Refiner
The Syntax Refiner catches and corrects superficial compilation errors: missing commas, wrong keywords (e.g. Lean3’s from vs. Lean4’s := by), misplaced brackets, etc. These syntax errors arise when a general-purpose LLM (such as o3-mini [39] or o4-mini [40]) has not been specialized on Lean4 syntax. By normalizing these common mistakes, this module ensures that subsequent stages operate on syntactically valid code.
@@ 3.2 Sorrifier
The Sorrifier patches any remaining compilation failures by inserting Lean’s sorry placeholder. We first parse the failed proof into a tree of nested proof-blocks (sub-lemmas as children). Then we compile the proof with Lean REPL [38], detect the offending line or block, and apply one of three repairs:
@@
    Block removal, if the entire sub-proof is malformed.
@@
    Insert sorry, if the block type-checks but leaves goals open.
We repeat this procedure until the file compiles without errors. At that point, every remaining sorry marks a sub-lemma to be proved in later stages of Apollo. This part of the pipeline guarantees that the proof successfully compiles in Lean with warnings of presence of sorry statements.
Each sorry block corresponds to a sub-problem that the LLM failed to prove. Such blocks may not type-check for various reasons, most commonly LLM hallucination. The feedback obtained via REPL lets us to automatically catch these hallucinations, insert a sorry placeholder, and remove invalid proof lines.
@@ 3.3 Auto Solver
The Auto Solver targets each sorry block in turn. It first invokes the Lean4’s hint to close the goal. If goals persist, it applies built-in solvers (nlinarith, ring, simp, etc.) wrapped in try to test combinations automatically. Blocks whose goals remain open stay marked with sorry. After applying Auto Solver, a proof may be complete with no sorry’s in which case Apollo has already succeeded in fixing the proof. Otherwise, the process can repeat recursively.
3.4 Recursive repair
In the case where a proof still has some remaining sorry’s after applying the Auto Solver, Apollo can consider each of the sorry’s as a new lemma, i.e., extract its local context (hypotheses, definitions, and any lemmas proved so far), and recursively try to prove the lemmas by prompting the LLM and repeating the whole process of verification, syntax refining, sorrifying, and applying the Auto Solver.
At each of these recursive iterations, a lemma may be completely proved, or it may end up with an incomplete proof with one or a few sorry’s. This allows the Apollo to make progress in proving the original theorem by breaking down the incomplete steps further and further until the LLM or Auto Solver can close the goals. This process can continue up to a user-specified recursion depth $r$.
@@ 3.5 Proof Assembler
Finally, the Proof Assembler splices all repaired blocks back into a single file and verifies that no sorry or admit (alias for "sorry") commands remain. If the proof still fails, the entire pipeline repeats (up to a user-specified recursion depth $r$), allowing further rounds of refinement and sub-proof generation.
Apollo’s staged approach: syntax cleaning, “sorrifying,” automated solving, and targeted LLM-driven repair, yields improvements in both proof-sample efficiency and final proof correctness.
@@ 4 Experiments
In this section, we present empirical results for Apollo on the MiniF2F dataset [8], which comprises of 488 formalized problems drawn from AIME, IMO, and AMC competitions. The benchmark is evenly split into validation and test sets (244 problems each); here we report results on the MiniF2F-test partition. To demonstrate Apollo’s effectiveness, we evaluate it on a range of state-of-the-art whole-proof generation models. All experiments use Lean v4.17.0 and run on eight NVIDIA A5000 GPUs with 128 CPU cores. We used @32@32 sampling during sub-proof generation unless stated otherwise. Baseline numbers are sourced from each model’s original publication works.
@@ 4.1 The effect of applying Apollo on top of the base models
Figure 3 shows the impact of Apollo on two state-of-the-art whole-proof generation models: Goedel Prover-SFT [31] and Kimina-Prover-Preview-Distill-7B [34]. Applying Apollo increases Goedel-Prover-SFT’s top accuracy from 64.7% to 65.6% while reducing its sampling budget by two orders of magnitude (from 25,600 generated samples to only a few hundred on average). Similarly, Kimina-Prover-Preview-Distill-7B achieves a new best Kimina-Prover-Preview-Distill-7B accuracy of 75.0% with roughly one-third of the previous sample budget. We report the exact numbers in Table 1.
@@
Note that Apollo’s sampling budget is not fixed: it depends on the recursion depth $r$ and the LLM’s ability to generate sub-lemmas. For each sub-lemma that Auto Solver fails to close, we invoke the LLM with a fixed top-32 sampling budget. Because each generated sub-lemma may spawn further sub-lemmas, the total sampling overhead grows recursively, making any single @K budget difficult to report. Sub-lemmas also tend to require far fewer tactics (and thus far fewer tokens) than the main theorem. A theorem might need 100 lines of proof while a sub-lemma might need only 1 line. Hence, sampling cost does not scale one-to-one. Therefore, to compare with other whole-proof generation approaches, we report the average number of proof samples needed to reach a fully verified solution.
Table 1: Table results comparing sampling cost, accuracy, and token usage before and after applying Apollo. “Cost” denotes the sampling budget $K$ for whole-proof generation and the recursion depth $r$ for Apollo. Column NN reports the average number of tokens generated per problem in "Chain-of-Thought" mode. Bold cells highlight the best accuracy. Results are shown on the miniF2F-test dataset for two models: Kimina-Prover-Preview-Distill-7B and Goedel-SFT.
Goedel-Prover-SFT Kimina-Prover-Preview-Distill-7B

### v4

@@ Abstract
Formal reasoning and automated theorem proving constitute a challenging subfield of machine learning, in which machines are tasked with proving mathematical theorems using formal languages like Lean. A formal verification system can check whether a formal proof is correct or not almost instantaneously, but generating a completely correct formal proof with large language models (LLMs) remains a formidable task. The usual approach in the literature is to prompt the LLM many times (up to several thousands) until one of the generated proofs passes the verification system. In this work, we present APOLLO (Automated PrOof repair via LLM and Lean cOllaboration), a modular, model‑agnostic agentic framework that combines the strengths of the Lean compiler with an LLM’s reasoning abilities to achieve better proof‐generation results at a low token and sampling budgets. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub‑lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top‑$K$ budget. The repaired sub‑proofs are recombined and reverified, iterating up to a user‑controlled maximum number of attempts. On the miniF2F benchmark, we establish a new state‑of‑the‑art accuracy of 84.9% among sub 8B‑parameter models (as of August 2025) while keeping the sampling budget below one hundred. Moreover, Apollo raises the state‑of‑the‑art accuracy for Goedel‑Prover‑SFT to 65.6% while cutting sample complexity from 25,600 to a few hundred. General‑purpose models (o3‑mini, o4‑mini) jump from 3–7% to over 40% accuracy. Our results demonstrate that targeted, compiler‑guided repair of LLM outputs yields dramatic gains in both efficiency and correctness, suggesting a general paradigm for scalable automated theorem proving. The codebase is available at https://github.com/aziksh-ospanov/APOLLO
@@ 1 Introduction
Formal reasoning has emerged as one of the most challenging fields of AI with recent achievements such as AlphaProof and AlphaGeometry doing well at the International Math Olympiad competing with humans yang2024formal ; alphaproof ; AlphaGeometryTrinh2024 . Formal reasoning relies both on AI models and a formal verification system that can automatically verify whether a mathematical proof is correct or not. In recent years, formal verification systems such as Lean lean4paper have facilitated a new form of doing mathematical research by enabling mathematicians to interact with the formal verification system and also with each other via the system, enabling larger numbers of mathematicians to collaborate with each other on a single project. As such, these formal verification systems are also called proof assistants as one can use them interactively to write a formal proof and instantaneously observe the current state of the proof and any possible errors or shortcomings in the proof generated by the compiler. Immediate access to the output of Lean compiler can help a mathematician to fix possible errors in the proof. At the same time, when a proof passes the compiler with no error, other collaborators do not need to verify the correctness of the proof. This type of collaboration is transforming the field of mathematics enabling large groups of collaborators to engage in large projects of mathematical research such as the Prime Number Theorem And More project kontorovich_tao_2024_primenumbertheoremand . Moreover, it has given rise to the digitization of mathematics buzzard2024mathematical .
In the AI front, large language models (LLMs) are improving at mathematical reasoning in natural language, and at the same time, they have shown a remarkable ability to learn the language of those formal verification systems and write mathematical proofs in formal language. This has led to the field of Automated Theorem Proving (ATP) where the standard approach is to prompt the LLM to generate a number of candidate proofs for a given theorem which will then be verified automatically using the formal verification system. Better models, better training sets, and better training methods has led to significant advances in this field tsoukalas2024putnambenchevaluatingneuraltheoremprovers ; zheng2022minif2fcrosssystembenchmarkformal ; azerbayev2023proofnetautoformalizingformallyproving ; li2024numinamath ; hu2025minictx .
@@
In this work, we address this issue by introducing Apollo, a fully automated system which uses a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with the reasoning abilities of LLMs. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub‑lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top‑$K$ budget. The high-level overview of Apollo appears at the bottom of Figure 1.
@@
  • We establish a new SOTA on miniF2F-test benchmark for medium-sized language models. (with parameter size 8B or less)
@@ 2 Related Works
Formal theorem proving systems. At their core, formal theorem‑proving (FTP) systems employ interactive theorem provers (ITPs) to verify mathematical results. In particular, Lean4 lean4paper is both a functional programming language and an ITP. Each theorem, lemma, conjecture, or definition must be checked by Lean’s trusted kernel, so the compiler’s verdict is binary: either the statement type‑checks (True) or it fails (False). This rigorous verification dramatically increases the reliability of formal mathematics; however, it also increases the complexity of proof development: authors must both comprehend the mathematical concepts and precisely encode them in the formal language. The usefulness of Lean is also shown to go beyond theorem proving, as Jiang et al. jiang2024leanreasoner showed that Lean can be adapted to natural language logical reasoning.
@@
Patching broken programs. Many prior works have explored using feedback to repair broken proofs. In software engineering, “repair” typically refers to program repair, i.e. fixing code that no longer runs programrepair . A common source of errors is a version mismatch that introduces bugs and prevents execution. For example, SED sed_repair uses a neural program‐generation pipeline that iteratively repairs initial generation attempts. Other notable approaches train specialized models to fix broken code based on compiler feedback (e.g. BIFI bifi , TFix tfix ) or leverage unsupervised learning over large bug corpora (e.g. BugLab buglab ).
@@
LLM collaboration with external experts. The use of expert feedback, whether from human users or from a proof assistant’s compiler, has proven effective for repairing broken proofs. Ringer ringer2009typedirected showcased automatic proof repair within the Coq proof assistant bertot2013interactive , enabling the correction of proofs in response to changes in underlying definitions. Jiang et al. jiang2022thor showed that leveraging automatic solvers with generative models yield better performance on formal math benchmarks. More recently, First et al. first2023baldurwholeproofgenerationrepair showed that incorporating compiler feedback into LLM proof generation significantly improves the model’s ability to correct errors and produce valid proofs. Another line of work song2023towards explored an idea of mathematician and LLM collaboration, where human experts are aided by LLMs at proof writing stage.
@@
Proof search methods. Numerous recent systems use machine learning to guide the search for formal proofs. One of the methods involves using large language models with structured search strategies, e.g. Best‑First Search (BFS) polu2022formalmathematicsstatementcurriculum ; wu2024internlm25stepproveradvancingautomatedtheorem ; xin2025bfsproverscalablebestfirsttree or Monte Carlo Tree Search (MCTS) lample2022hypertree ; wang2023dt-solver ; alphaproof . While tree‑search methods reliably steer a model toward valid proofs, they result in high inference costs and often explore many suboptimal paths before success. Another line of work leverages retrieval based systems to provide context for proof generators with potentially useful lemmas. One such example is ReProver yang_leandojo_2023 that augments Lean proof generation by retrieving relevant lemmas from a proof corpus and feeding them to an LLM, enabling the model to reuse existing proofs.
@@
Whole proof generation. The use of standalone LLMs for theorem proving has emerged as a major research direction in automated theorem proving. One of the earliest works presented GPT-f polu2020generativelanguagemodelingautomated , a transformer-based model for theorem proving that established the LLMs ability in formal reasoning. As training frameworks and methods advance, the community has produced numerous models that generate complete proofs without external search or tree‑search algorithms. Recent work shows that both supervised models xin2024deepseekproverv15harnessingproofassistant ; lin2025goedelproverfrontiermodelopensource and those trained via reinforcement learning xin2024deepseekproverv15harnessingproofassistant ; zhang2025leanabellproverposttrainingscalingformal ; dong2025stp ; kimina_prover_2025 achieve competitive performance on formal‑math benchmarks. However, whole‑proof generation remains vulnerable to hallucinations that can cause compilation failures even for proofs with correct reasoning chains.
@@
Informal Chain-of-Thought in Formal Mathematics. Several recent works demonstrate that interleaving informal “chain‑of‑thought” (CoT) reasoning with formal proof steps substantially improves whole‑proof generation. For example, developers insert natural‑language mathematical commentary between tactics, yielding large accuracy gains lin2025goedelproverfrontiermodelopensource ; xin2024deepseekproverv15harnessingproofassistant ; azerbayev2024llemma . Wang, et al. kimina_prover_2025 further show that special “thinking” tokens let the model plan its strategy and self‑verify intermediate results. The “Draft, Sketch, and Prove” (DSP) jiang2023draftsketchproveguiding and LEGO-Prover wang2024legoprover pipelines rely on an informal proof sketch before attempting to generate the formal proof. Collectively, these studies indicate that providing informal mathematical context, whether as comments, sketches, or dedicated tokens, guides LLMs toward correct proofs and significantly boosts proof synthesis accuracy.
Figure 2: Overview of the Apollo framework.
@@ 3 Our Approach
Our compiler of choice is the Lean REPL leanrepl , a “Read–Eval–Print–Loop” for Lean4. It provides a complete list of compilation errors and warnings, each with a detailed description and the location (for example, the line number) where it occurred. Errors can include syntax mistakes, nonexistent tactics, unused tactics, and more. We aggregate this information to drive our agent’s dynamic code repair and sub‑problem extraction. The pseudo-code for our framework can be found in Algorithm 1 in the Appendix.
@@ 3.1 Syntax Refiner
The Syntax Refiner catches and corrects superficial compilation errors: missing commas, wrong keywords (e.g. Lean3’s from vs. Lean4’s := by), misplaced brackets, etc. These syntax errors arise when a general‑purpose LLM (such as o3‑mini openai-o3-mini-2025 or o4‑mini openai-o4-mini-2025 ) has not been specialized on Lean4 syntax. By normalizing these common mistakes, this module ensures that subsequent stages operate on syntactically valid code. It is important to note that this module is primarily applicable to general purpose models not explicitly trained for theorem proving in Lean. In contrast, Syntax Refiner usually does not get invoked for proofs generated by LLMs trained for formal theorem proving.
@@ 3.2 Sorrifier
The Sorrifier patches any remaining compilation failures by inserting Lean’s sorry placeholder. We first parse the failed proof into a tree of nested proof‑blocks (sub‑lemmas as children). Then we compile the proof with Lean REPL leanrepl , detect the offending line or block, and apply one of three repairs:
@@
    Block removal, if the entire sub‑proof is malformed.
@@
    Insert sorry, if the block compiles but leaves unsolved goals open.
We repeat this procedure until the file compiles without errors. At that point, every remaining sorry marks a sub‑lemma to be proved in later stages of Apollo. This part of the pipeline guarantees that the proof successfully compiles in Lean with warnings of presence of sorry statements.
Each sorry block corresponds to a sub‑problem that the LLM failed to prove. Such blocks may not type‑check for various reasons, most commonly LLM hallucination. The feedback obtained via REPL lets us to automatically catch these hallucinations, insert a sorry placeholder, and remove invalid proof lines.
@@ 3.3 Auto Solver
The Auto Solver targets each sorry block in turn. It first invokes the Lean4’s hint to close the goal. If goals persist, it applies built‑in solvers (nlinarith, ring, simp, etc.) wrapped in try to test combinations automatically. Blocks whose goals remain open stay marked with sorry. After applying Auto Solver, a proof may be complete with no sorry’s in which case Apollo has already succeeded in fixing the proof. Otherwise, the process can repeat recursively.
3.4 Recursive reasoning and repair
In the case where a proof still has some remaining sorry statements after applying the Auto Solver, Apollo can consider each as a new lemma, i.e., extract its local context (hypotheses, definitions, and any lemmas proved so far), and recursively try to prove the lemmas by prompting the LLM and repeating the whole process of verification, syntax refining, sorrifying, and applying the Auto Solver.
At each of these recursive iterations, a lemma may be completely proved, or it may end up with an incomplete proof with one or a few sorry blocks. This allows the Apollo to make progress in proving the original theorem by breaking down the incomplete steps further and further until the LLM or Auto Solver can close the goals. This process can continue up to a user‑specified recursion depth $r$.
@@ 3.5 Proof Assembler
Finally, the Proof Assembler splices all repaired blocks back into a single file and verifies that no sorry or admit (alias for "sorry") commands remain. If the proof still fails, the entire pipeline repeats (up to a user‑specified recursion depth $r$), allowing further rounds of refinement and sub‑proof generation.
Apollo’s staged approach: syntax cleaning, “sorrifying,” automated solving, and targeted LLM‐driven repair, yields improvements in both proof‐sample efficiency and final proof correctness.
@@ 4 Experiments
In this section, we present empirical results for Apollo on the miniF2F dataset zheng2022minif2fcrosssystembenchmarkformal , which comprises of 488 formalized problems drawn from AIME, IMO, and AMC competitions. The benchmark is evenly split into validation and test sets (244 problems each); here we report results on the miniF2F‑test partition. To demonstrate Apollo’s effectiveness, we evaluate it on a range of state‑of‑the‑art whole‑proof generation models. All experiments use Lean v4.17.0 and run on eight NVIDIA A5000 GPUs with 128 CPU cores. We used @32@32 sampling during sub-proof generation unless stated otherwise. Baseline numbers are sourced from each model’s original publication works.
@@ 4.1 The effect of applying Apollo on top of the base models
Figure 3 shows the impact of Apollo on two state‑of‑the‑art whole‑proof generation models: Goedel Prover‑SFT lin2025goedelproverfrontiermodelopensource and Kimina‑Prover‑Preview‑Distill‑7B kimina_prover_2025 . Applying Apollo increases Goedel‑Prover‑SFT’s top accuracy from 64.7% to 65.6% while reducing its sampling budget by two orders of magnitude (from 25,600 generated samples to only a few hundred on average). Similarly, Kimina‑Prover‑Preview‑Distill‑7B achieves a new best Kimina-Prover-Preview-Distill-7B accuracy of 75.0% with roughly one‑third of the previous sample budget. We report the exact numbers in Table 1.
@@
Note that Apollo’s sampling budget is not fixed: it depends on the recursion depth $r$ and the LLM’s ability to generate sub‑lemmas. For each sub‑lemma that Auto Solver fails to close, we invoke the LLM with a fixed top‑32 sampling budget. Because each generated sub‑lemma may spawn further sub‑lemmas, the total sampling overhead grows recursively, making any single @K budget difficult to report. Sub‑lemmas also tend to require far fewer tactics (and thus far fewer tokens) than the main theorem. A theorem might need 100 lines of proof while a sub-lemma might need only 1 line. Hence, sampling cost does not scale one‑to‑one. Therefore, to compare with other whole-proof generation approaches, we report the average number of proof samples and token budgets. We present more in-depth analysis of inference budgets in Section B of the Appendix.
Table 1: Table results comparing sampling cost, accuracy, and token usage before and after applying Apollo. “Cost” denotes the sampling budget $K$ for whole‑proof generation and the recursion depth $r$ for Apollo. Column NN reports the average number of tokens generated per problem in "Chain‑of‑Thought" mode. Bold cells highlight the best accuracy. Results are shown on the miniF2F‑test dataset for two models: Kimina‑Prover‑Preview‑Distill‑7B and Goedel‑SFT.
Goedel‑Prover‑SFT Kimina‑Prover‑Preview‑Distill‑7B